OpenStudy (anonymous):

the data below gives the number of bacteria, in millions, found in a certain culture. CLICKS 2 SEE PLEASE.

OpenStudy (anonymous):

Time (hours) -- 0 1 2 3 4

OpenStudy (anonymous):

Bacteria ---- 5 8 15 26 48 , consecutively

OpenStudy (anonymous):

a. find an exponential function that models the data . B... write the equation part a in terms of base e. C. use the model to estimate the doubling time for the culture.

OpenStudy (tkhunny):

If we have to use an Expoential Model, Think Ratios! 8/5 = 1.6 15/8 = 1.875 26/15 = 1.7333 48/26 = 1.8462 Okay, it's not perfect. How do you suppose we should build the model? Just guess? Least Squares?

OpenStudy (anonymous):

using a formula .

OpenStudy (anonymous):

a specific formula ?

OpenStudy (tkhunny):

Well, yes, we are commanded to use an exponential formula, but how? Which one? I'm going to guess you're not quite ready for Least Squares. Exponential formulas can be of all sorts: \(y=Ae^{bx}\) \(y=A+Be^{cx}\) \(y = \dfrac{A}{B + Ce^{dx}}\) Shall we just pick the simplest?

OpenStudy (tkhunny):

Sorry, the preview isn't working. It's a little tricky to code the LaTeX to get math expressions.

OpenStudy (anonymous):

the simplest, the 1st one ?

OpenStudy (tkhunny):

Fine. There is still an important choice to make. Since exponentials have CONSTANT ratios, and our data DON'T, what shall we do with that? Any ideas?

OpenStudy (anonymous):

make constant ratios out of our data?

OpenStudy (tkhunny):

Well, we can't just make it be constant when it isn't. The task is to pick a constant ration that approximates the data pretty well. Let's just use the first and last values and see how close we get in the middle. \(\left(\dfrac{48}{5}\right)^{1/4} = 1.76022\) See how this values matches NONE of the ratios we first calculated but does sort of hang around them all?

OpenStudy (anonymous):

yes .

OpenStudy (tkhunny):

Well, then let's just go with it. \(B(t) = 5\cdot 1.76022^{t}\) Try it out. You shoudl be in the neighborhoos for t = 1, 2, and 3, but right on for t = 0 and 4. Lastly, let's review to see if we have fulfilled the requirements of the original problem statement.

OpenStudy (anonymous):

you want me to give you the values of everything once i place 1,2, and 3, in for t?

OpenStudy (tkhunny):

You don't have to give them to me. We're trying to convince you that it's a rational model. You need to believe!

OpenStudy (anonymous):

for 2 i got 15.491872242

OpenStudy (anonymous):

for 3 i got 27.26910335781324

OpenStudy (tkhunny):

I think you have it. What say you? Are they looking like they are in the neighborhood of the provided values?

OpenStudy (anonymous):

yes .

OpenStudy (tkhunny):

Okay, I think we're done with Part A. How shall we do Part B?

OpenStudy (anonymous):

write an equation from what we just got in base of e terms?

OpenStudy (tkhunny):

That's it. We have to solve this: \(1.76022 = e^{b}\). It is a delightful little logarithm problem. Here's your chance! Use logarithms to sovle for 'b'.

OpenStudy (anonymous):

ln (e^ 1.76022) = ln e^b ?

OpenStudy (tkhunny):

No, there was no 'e' on the left side.